OpenStudy (anonymous):

Sue and Fred have a combined age of 42 years. If Sue's age is 1/2 of Fred's, how old is Fred in years? 16 20 23 28 35

OpenStudy (anonymous):

sue's age = S fred's age = F S + F = 42 S = (1/2)F substitute for S (1/2)F + (1)F = 42 (3/2)F = 42 multiply both sides by 2/3 (2/3)(3/2)F = 42(2/3); F = 42(2/3) = 14(2) = 28 S + 28 = 42 S = 42 - 28 = 14 Look back at the problem to see what they are asking. Fred is 28